# Neural Graph Pattern Machine

### v1

@@ 4.2 Out-of-Distribution Generalization
We evaluate the model robustness under distribution shifts between training and testing sets. Under the setting, the model is trained on a source graph and evaluated on a target graph, with a 20/80 val/test split. We conduct experiments on citation networks, ACM and DBLP (using accuracy as the metric), as well as social networks Twitch (using AUC as the metric). The Twitch dataset consists of six graphs (DE, EN, ES, FR, PT, RU), where the model is trained on DE and evaluated on the remaining graphs. For baselines, in addition to the methods used in previous experiments, we include domain adaptation and OOD generalization baselines such as DANN, SR-GNN, StruRW, and SSReg. The experimental results are summarized in Table 4. We observe that standard graph learning methods struggle in this setting, highlighting their limited robustness to OOD testing. In contrast, our GPM outperforms existing OOD-specific methods in settings such as A $\to$ D and Twitch, demonstrating superior robustness to OOD challenges. This can be attributed to GPM’s pattern learning ability that potentially identifies shared patterns between source and target graphs, whereas message passing is sensitive to subtle structural changes (Wang et al., 2024d). Furthermore, the performance of GPM is enhanced when combined with OOD techniques such as DANN and SSReg, achieving significant improvements across all settings, particularly in D $\to$ A.
@@ 4.3 Scalability
Distributed Training. Transformers have demonstrated remarkable efficiency in distributed training due to the parallelization of self-attention mechanism (Shoeybi et al., 2019). In contrast, message passing GNNs are less efficient for distributed training, as their iterative message passing introduces sequential dependencies and incurs significant communication overhead when computational nodes are distributed across machines. By leveraging a transformer-based architecture, GPM achieves superior efficiency in distributed training compared to message passing GNNs on Products, as shown in Figure 4 (Bottom). Further details and discussions are provided in Appendix G.1.
@@ 4.4 Ablation Study
Impact of Model Components. Table 5 presents the model component ablation study results. Both positional embeddings (PE) and anonymous paths (AP) contribute to capturing topological information, with PE encoding relative node positions and AP characterizing pattern structures. Empirically, AP has a greater impact than PE, suggesting that the model prioritizes understanding pattern structures over node locations. For the semantic path (SP) encoder, the Transformer achieves the best due to its ability to adaptively model both localized and long-range dependencies. More results and discussions are provided in Appendix G.2.
Impact of Learning Strategies. Multi-scale training enhances the average performance from 70.72 to 72.34. Test-time augmentation significantly accelerates training without sacrificing performance, with an average score of 72.34 compared to 72.43. Details are provided in Appendix G.3.
@@ 5 Conclusion
Boarder Impact
@@ Appendix A Limitations of Message Passing and Recent Advances
The message passing paradigm (Kipf & Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018; Wang et al., 2023, 2024b, 2024a; Zhang et al., 2024b, c; Wang et al., 2024c) in GNNs has well-documented limitations, including restricted expressiveness, over-smoothing, over-squashing, and an inability to effectively model long-range dependencies. Given our focus on learning graph patterns, this discussion emphasizes the issues of expressiveness and notable advancements. Pioneering work by Xu et al. (2019) revealed that the expressive power of GNNs is fundamentally bounded by the 1-WL isomorphism test. Building on this, Corso et al. (2020) demonstrated that no GNN employing a single aggregation function can achieve the expressiveness of the 1-WL test when the neighborhood multiset has uncountable support. Consequently, this constraint renders GNNs incapable of identifying critical graph structures, such as stars, conjoint cycles, and $k$-cliques (Chen et al., 2020; Garg et al., 2020; Zhang et al., 2024a).
@@
An alternative approach leverages random walk kernels to guide the message passing process, constraining interactions to a limited range of nodes (Jin et al., 2022; Martinkus et al., 2023; Tönshoff et al., 2023; Wang & Cho, 2024). Notably, Tönshoff et al. (2023) demonstrated that random walk-based methods can capture both small substructures and long-range dependencies, while Wang & Cho (2024) showed that sufficiently long random walks can distinguish non-isomorphic graphs. Despite their strengths, these methods tend to emphasize long-range dependencies at the expense of localized information (Tönshoff et al., 2023) and lack interpretability regarding the specific graph knowledge being learned.
@@ B.4 Architectures in Model Scaling Analysis
# Transformer Layers 2 2 2 3 3
# Number of Heads 8 24 32 48 112
# Hidden Dimension 512 1536 2048 3072 7168
@@ Appendix C Proof
(1) The simplified case assumes $n=1$ with a sufficiently large walk length $l$, whereas GPM operates with $n>1$. Since random walks can start from any node in the graph, an $l$-length random walk can be split into $k$ segments ($k$ is large enough), each of length $l/k$. Each sub-walk can be encoded individually and later combined to form the final embedding. (2) GPM employs neural networks as encoders, which are inherently universal and injective, satisfying the requirements of the mappings $\rho_{s}$ and $\rho_{a}$ used in the simplified case. (3) Finally, while the simplified case uses a mean aggregator over encoded patterns, GPM adopts a transformer architecture for aggregation. The self-attention mechanism in transformers generalizes the mean aggregator, which can be seen as a special case of the transformer.
@@ Appendix D Complexity
Appendix E Limitations
The prediction performance of GPM heavily depends on the quality of the sampled patterns. In this work, we adopt a random sampling strategy, aiming to sample as many patterns as possible to construct a comprehensive pattern set. However, this approach may increase resource consumption during model training and hyperparameter tuning. Developing an adaptive sampling strategy tailored to specific downstream tasks or designing a unified pattern vocabulary could mitigate this issue. Furthermore, the current implementation of GPM is limited to supervised tasks. Future extensions could include unsupervised learning, integration with LLMs, or adaptations for complex graph types, such as heterogeneous graphs.
Appendix F How Does GPM Surpass Message Passing?
F.1 Empirical Effectiveness
@@
F.2 Tackling Over-Squashing
@@
Appendix G Additional Experimental Results and Discussions
G.1 Additional Discussion on Distribution Training
@@
These findings highlight the advantages of GPM, which using tranformer as backbone, over message-passing GNNs for distributed training, particularly in scenarios with a large number of devices.
G.2 Model Component Ablation
@@
G.3 Additional Discussion on Learning Strategies
@@
Appendix H GPM Automatically Learns Data Dependencies
@@
Appendix I Model Interpretation on Computers

### v2

@@ 4.1 Applicability to Graph-based Tasks
To evaluate the effectiveness of our method on all graph predictive tasks (node, link, and graph), we conduct extensive experiments such as node classification, link prediction, graph classification, and graph regression.
@@ 4.2 Out-of-Distribution Generalization
We evaluate the model robustness under distribution shifts between training and testing sets. Under the setting, the model is trained on a source graph and evaluated on a target graph, with a 20/80 val/test split. We conduct experiments on citation networks, ACM and DBLP (using accuracy as the metric), as well as social networks Twitch (using AUC as the metric). The Twitch dataset consists of six graphs (DE, EN, ES, FR, PT, RU), where the model is trained on DE and evaluated on the remaining graphs. For baselines, in addition to the methods used in previous experiments, we include domain adaptation and OOD generalization baselines such as DANN, SR-GNN, StruRW, and SSReg. The experimental results are summarized in Table 4. We observe that standard graph learning methods struggle in this setting, highlighting their limited robustness to OOD testing. In contrast, our GPM outperforms existing OOD-specific methods in settings such as A $\to$ D and Twitch, demonstrating superior robustness to OOD challenges. This can be attributed to GPM’s pattern learning ability that potentially identifies shared patterns between source and target graphs, whereas message passing is sensitive to subtle structural changes (Wang et al., 2024c). Furthermore, the performance of GPM is enhanced when combined with OOD techniques such as DANN and SSReg, achieving significant improvements across all settings, particularly in D $\to$ A.
@@ 4.3 Scalability
Distributed Training. Transformers have demonstrated remarkable efficiency in distributed training due to the parallelization of self-attention mechanism (Shoeybi et al., 2019). In contrast, message passing GNNs are less efficient for distributed training, as their iterative message passing introduces sequential dependencies and incurs significant communication overhead when computational nodes are distributed across machines. By leveraging a transformer-based architecture, GPM achieves superior efficiency in distributed training compared to message passing GNNs on Products, as shown in Figure 4 (Bottom). Further details and discussions are provided in Appendix F.1.
@@ 4.4 Ablation Study
Impact of Model Components. Table 5 presents the model component ablation study results. Both positional embeddings (PE) and anonymous paths (AP) contribute to capturing topological information, with PE encoding relative node positions and AP characterizing pattern structures. Empirically, AP has a greater impact than PE, suggesting that the model prioritizes understanding pattern structures over node locations. For the semantic path (SP) encoder, the Transformer achieves the best due to its ability to adaptively model both localized and long-range dependencies. More results and discussions are provided in Appendix F.2.
Impact of Training Tricks. Both multi-scale training and test-time augmentation contribute to improved performance. Specifically, multi-scale training increases average accuracy from 70.72 to 72.34 by incorporating hierarchical substructure knowledge. Test-time augmentation uses fewer substructure patterns during training (e.g., 16 at training vs. 128 at inference), which significantly reduces training cost with negligible performance drop (72.43 to 72.34). Full details are provided in Appendix F.3.
@@
GPM Automatically Learns Data Dependencies. Leveraging the transformer architecture, GPM automatically identifies dominant patterns. Experiments on the Products (Figure 5) demonstrate that unbiased random walk sampling ($p=1,q=1$) achieves the best results, over localized ($p=0.1,q=10$) and long-range ($p=10,q=0.1$) sampling, by allowing the model to autonomously balance localized and long-range dependencies. Detailed discussions are provided in Appendix G.
4.5 Model Interpretation
@@ 5 Conclusion
Limitations and Future Works. The prediction performance of GPM heavily depends on the quality of the sampled patterns. In this work, we adopt a random sampling strategy, aiming to sample as many patterns as possible to construct a comprehensive pattern set. However, this approach may increase resource consumption during model training and hyperparameter tuning. Developing an adaptive sampling strategy tailored to specific downstream tasks or designing a unified pattern vocabulary (Wang et al., 2024a, b) could mitigate this issue. Furthermore, the current implementation of GPM is limited to supervised tasks. Future extensions could include unsupervised learning (He et al., 2022), integration with LLMs (Yuan et al., 2021; Liu et al., 2024), incorporating external knowledge (Ni et al., 2025), or adaptations for complex graph types, such as heterogeneous graphs (Wang et al., 2023).
Acknowledgments
This work was partially supported by the NSF under grants IIS-2321504, IIS-2334193, IIS-2340346, IIS-2217239, CNS-2426514, and CMMI-2146076, and Notre Dame Strategic Framework Research Grant (2025). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
Impact Statement
@@ Appendix A Limitations of Message Passing and Recent Advances
The message passing paradigm (Kipf & Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018) in GNNs has well-documented limitations, including restricted expressiveness, over-smoothing, over-squashing, and an inability to effectively model long-range dependencies. Given our focus on learning graph patterns, this discussion emphasizes the issues of expressiveness and notable advancements. Pioneering work by Xu et al. (2019) revealed that the expressive power of GNNs is fundamentally bounded by the 1-WL isomorphism test. Building on this, Corso et al. (2020) demonstrated that no GNN employing a single aggregation function can achieve the expressiveness of the 1-WL test when the neighborhood multiset has uncountable support. Consequently, this constraint renders GNNs incapable of identifying critical graph structures, such as stars, conjoint cycles, and $k$-cliques (Chen et al., 2020; Garg et al., 2020; Zhang et al., 2024a).
@@
An alternative approach leverages random walk kernels to guide the message passing process, constraining interactions to a limited range of nodes (Jin et al., 2022; Martinkus et al., 2023; Tönshoff et al., 2023; Wang & Cho, 2024; Chen et al., 2025). Notably, Tönshoff et al. (2023) demonstrated that random walk-based methods can capture both small substructures and long-range dependencies, while Wang & Cho (2024) showed that sufficiently long random walks can distinguish non-isomorphic graphs. Moreover, these random walk approaches are theoretically more expressive than message passing (Zhang et al., 2019; Fan et al., 2022; Welke et al., 2023; Michel et al., 2023; Graziani et al., 2024). Despite their strengths, these methods tend to emphasize long-range dependencies at the expense of localized information (Tönshoff et al., 2023) and lack interpretability regarding the specific graph knowledge being learned.
@@ B.3 Model Configurations
Table 6: Hyper-parameter settings on predictive tasks.
Products Computer Arxiv WikiCS CoraFull Deezer Blog Flickr Flickr-S
Task Node Node Node Node Node Node Node Node Node
Learning Rate 0.01 0.01 0.01 0.01 0.01 0.01 0.001 0.01 0.005
Dropout 0.3 0.1 0.1 0.5 0.1 0.1 0.1 0.5 0.3
Decay 0 0.01 0 0.01 0 0 0.01 0.01 0.01
Batch Size 256 256 256 256 256 256 256 256 256
PE Type Lap Lap Lap Lap Lap Lap Lap Lap Lap
PE Dim 4 8 4 8 8 16 16 8 4
AP Encoder MEAN GRU GRU GRU GRU GRU GRU MEAN GRU
$\lambda$ 0.5 0.5 1 1 1 0.5 0.1 0.5 0.5
Cora Pubmed ogbl-Collab IMDB-B COLLAB Reddit-M5K Reddit-M12K ZINC ZINC-Full
Task Link Link Link Graph Graph Graph Graph Graph Graph
Learning Rate 0.001 0.01 0.01 0.001 0.01 0.001 0.005 0.01 0.01
Dropout 0.3 0.3 0.1 0.1 0.1 0.1 0.1 0.1 0.1
Decay 0 0.01 0 0 0 0 0.01 0 0
Batch Size 256 256 256 1024 1024 1024 256 1024 1024
PE Type Lap Lap Lap RW RW RW RW RW RW
PE Dim 4 4 4 8 4 8 8 8 8
AP Encoder GRU GRU GRU MEAN MEAN MEAN GRU GRU GRU
$\lambda$ 0.5 1 0.5 0.1 0.1 0.1 0.1 1 1
@@ B.4 Architectures in Model Scaling Analysis
# Transformer Layers 1 1 3 3 3
# Number of Heads 4 4 4 8 16
# Hidden Dimension 64 256 256 512 1024
@@ Appendix C Proof
(1) The simplified case assumes $n=1$ with a sufficiently large walk length $l$, whereas GPM operates with $n>1$. Since random walks can start from any node in the graph, an $l$-length random walk can be split into $k$ segments ($k$ is large enough), each of length $l/k$. Each sub-walk can be encoded individually and later combined to approximately form the final embedding. Note that We did not mean to suggest that long walks can be fully reconstructed from shorter ones, especially since anonymous walks cannot preserve node identity across segments. Rather, we intended to describe an approximate strategy, where long walks are segmented into shorter sub-walks, each encoded independently. This design allows the transformer to aggregate distributed long-range information across these sub-patterns. (2) GPM employs neural networks as encoders, which are inherently universal and injective, satisfying the requirements of the mappings $\rho_{s}$ and $\rho_{a}$ used in the simplified case. (3) Finally, while the simplified case uses a mean aggregator over encoded patterns, GPM adopts a transformer architecture for aggregation. The self-attention mechanism in transformers generalizes the mean aggregator, which can be seen as a special case of the transformer.
@@ Appendix D Complexity
Appendix E How Does GPM Surpass Message Passing?
E.1 Empirical Effectiveness
@@
E.2 Tackling Over-Squashing
@@
Appendix F Additional Experimental Results and Discussions
F.1 Additional Discussion on Distribution Training
@@
These findings highlight the advantages of GPM, which using tranformer as backbone, over message passing GNNs for distributed training, particularly in scenarios with a large number of devices.
F.2 Model Component Ablation
@@
F.3 Additional Discussion on Learning Strategies
@@
Appendix G GPM Automatically Learns Data Dependencies
@@
Appendix H Model Interpretation on Computers